force of 16 N acts tangent to the edge of a solid cylinder weighing 18 N and which has a

diameter of 76 cm. Determine the kinetic energy of the cylinder after 15 s.


1
Expert's answer
2021-03-12T07:13:12-0500

Rotational kinetic energy of the cylinder can be found as follows:


"KE=\dfrac{1}{2}I\omega^2."

The moment of inertia of the cylinder can be found as follows:


"I=\dfrac{1}{2}mr^2=\dfrac{1}{2}\cdot\dfrac{18\ N}{9.8\ \dfrac{m}{s^2}}\cdot(0.38\ m)^2=0.133\ kg\cdot m^2."

Let's find the angular acceleration of the cylinder:


"\tau=I\alpha,""\alpha=\dfrac{\tau}{I}=\dfrac{16\ N}{0.133\ kg\cdot m^2}=120.3\ \dfrac{rad}{s^2}."

Let's assume that cylinder begins to rotate from rest. Then, we can find its angular velocity after 15 seconds as follows:


"\omega=\omega_0+\alpha t=0+120.3\ \dfrac{rad}{s^2}\cdot15\ s=1804.5\ \dfrac{rad}{s}."

Finally, we can determine the kinetic energy of the cylinder after 15 s


"KE=\dfrac{1}{2}\cdot0.133\ kg\cdot m^2\cdot(1804.5\ \dfrac{rad}{s})^2=2.16\cdot10^5\ J."